OpenStudy (anonymous):

Find the coordinates of any points of intersection. Given: f(x)=x^2-5x+2 g(x)=3-2x

OpenStudy (anonymous):

First, set them equal to each other. \[x^2 -5x + 2 = 3 - 2x\] Then, get everything on one side to set it equal to zero. \[x^2 - 3x - 1\] Factor using the quadratic formula, you will get x = \[(3 + \sqrt{13}) \div 2, (3 - \sqrt{13}) \div 2\] Plug these x values into each equation to check (unfortunately, I don't have a good calculator on me right now, but the rest is just calculation. Plug in the x values to get the y values, and use them as your points.